MATH 208  SYLLABUS (MWRF)  SPRING 1999

TEXT:  Multivariable Calculus, McCallum, et al., 1st ed., Wiley (1997).
CALCULATOR:  A graphing calculator will be useful for this course. A TI-85 or TI-86 is recommended.

DAILY WORK: The exercises suggested below represent a minimal assignment. Some students may have to work additional exercises from the text to attain sufficient mastery of the material.

PREREQUISITES: Math 208 extends the concepts and techniques of Math 106 and Math 107. Students who are not thoroughly competent at this material will need to budget extra study time for review of 1-variable calculus.

NOTES: The pace of the course, order of topics, number of exams, projects, quizzes, and number of assignments may be varied at the discretion of your instructor. Your grade in the course will be determined by your instructor, who will describe his/her grading policy in detail.

EXAMS: Examinations may, in part, test whether the student can apply concepts learned in the course to new situations. Thus problems appearing on the exams may be different from exercises in the text.

FINAL EXAM: All Math 208 students are required to take a common comprehensive final examination covering all the topics listed on this syllabus, The final exam is scheduled for 6:00–8:00 p.m., MONDAY, MAY 3, 1999, in a room to be announced during the last week of class. You must arrange your personal and work schedules to allow you to take the exam at this scheduled time. Students with conflicting exam schedules may be permitted to take an alternate final exam AFTER the regularly scheduled exam.

CLASS DATE  SECTION(S)  EXERCISES
1  Jan. 11  Ch 11  p.56: 1–6, 11–15
2  Jan. 13  Ch 11  p.56: 16–19, 21, 23
3  Jan. 14  Ch 11
4  Jan. 15  12.1  p.66: 2, 3, 8, 11, 13, 17, 20, 23, 26, 28, 29
5  Jan. 20  12.2  p.74: 6, 8–10, 13, 15
6  Jan. 21  12.3  p.82: 4, 6–8, 11, 13, 17, 21, 24, 34
7  Jan. 22  12.4  p.90: 3, 7, 10, 13–16

Friday, January 22, is the last day to withdraw from the course and not have it appear on your transcript.

8  Jan. 25  Review
9  Jan. 27  13.1  p.102: 2, 3, 6, 8–10, 12, 15–17
10  Jan. 28  13.2  p.108: 1, 3, 7, 10, 12, 19, 21–23, 26, 28, 34, 35, 37, 38
11  Jan. 29  13.3  p.115: 3, 5, 10, 11, 13, 15, 19, 20, 22
12  Feb. 1  13.4  p.124: 2, 3, 5, 11, 13, 15, 20, 25, 28–30, 32, 36
13  Feb. 3  13.5  p.133: 2, 5, 8, 10–12, 14, 17, 18
14  Feb. 4  13.6  p.140: 1, 3, 5, 7, 9, 11, 13, 15, 17, 18, 25
15  Feb. 5  Review
16  Feb. 8  EXAM I
17  Feb. 10  13.7  p.145: 5–9, 11, 15, 17
18  Feb. 11  13.9  p.159: 1, 2, 4, 7, 10, 12
19  Feb. 12  11.7  p.55: 1, 2, 4, 5, 7, 9, 12, 14, 15
20  Feb. 15  13.10  p.167: 2(b–e), 3(b–e)
21  Feb. 17  Review
22  Feb. 18  14.1  p.183: 1, 2, 5–8, 13, 14, 16, 17
23  Feb. 19  14.2  p.192: 2, 3, 11, 17, 20

MATH 208  SYLLABUS  SPRING 1999
CLASS DATE  SECTION(S)  EXERCISES
24  Feb. 22  14.2  p.192: 5, 14, 21 (use a calculator) 23
25  Feb. 24  14.3  p.205: 1, 4, 5, 12, 14, 22, 23
26  Feb. 25  Review
27  Feb. 26  15.1  p.222: 1, 3, 4, 10, 11 (just do Nebraska)
28  Mar. 1  15.2  p.233: 1, 3, 5, 6, 9, 13–15, 17, 25
29  Mar. 3  15.3  p.238: 1, 5, 7, 9, 11, 13, 15, 16
30  Mar. 4  15.5  p.246: 1, 2, 5, 7, 9, 12, 13, 15, 16, 19
31  Mar. 5  15.6  p.254: 1–6, 11

Friday, March 5, is the last day to change your grade option to or from “Pass/No Pass”.

32  Mar. 8  15.6  p.254: 15, 16, 18, 21
33  Mar. 10  Review
34  Mar. 11  16.1  p.280: 1, 3, 6, 13, 15, 17, 19, 23, 25, 29, 31, 33
35  Mar. 12  16.2  p.291: 3, 5, 7, 13, 17, 23, 28, 29

The PROJECT will be assigned on March 12 and is due on April 7. Spring Break!

36  Mar. 22  Review
37  Mar. 24  EXAM II
38  Mar. 25  17.1  p.331: 5, 7
39  Mar. 26  17.2  p.337: 3, 5, 6
40  Mar. 29  18.1  p.348: 1, 3, 7, 9, 13, 15, 27
41  Mar. 31  18.2  p.356: 3, 4, 6, 7, 10–13
42  Apr. 1  Review
43  Apr. 2  18.3  p.364: 2–4, 8–11, 14–19
44  Apr. 5  18.4  p.375: 3, 6–9, 11, 13, 16
45  Apr. 7  Review
46  Apr. 8  19.1  p.392: 1, 3, 5, 7, 9, 11, 21, 22
47  Apr. 9  19.2  p. 403: 1, 3, 5, 8–10, 13, 15

Friday, April 9, is the last day to withdraw from the course and receive a grade of W.

48  Apr. 12  Review
49  Apr. 14  20.1  p.417: 5, 7, 8, 11, 17–19
50  Apr. 15  20.2  p.425: 1, 3, 6, 7, 9, 11, 12
51  Apr. 16  Review
52  Apr. 19  20.3  p.433: 2, 7, 11, 16, 17, 19–22
53  Apr. 21  20.4  p.440: 3, 5, 8, 15
54  Apr. 22  Review
55  Apr. 23  EXAM III
56  Apr. 26  Review
57  Apr. 28  Review
58  Apr. 29  Review
59  Apr. 30  Review

FINAL EXAM: Monday, May 3, 6:00–8:00 p.m.

Department Grading Appeals Policy: The Department of Mathematics and Statistics does not tolerate discrimination or harassment on the basis of race, gender, religion, or sexual orientation. If you believe you have been subject to such discrimination or harassment, in this or any other math course, please contact the department. If, for this or any other reason, you believe your grade was assigned incorrectly or capriciously, appeals may be made to (in order) the instructor, the department chair, the departmental grading appeals committee, the college grading appeals committee, and the university grading appeals committee.

NOTE: If this is your first college mathematics course, then you may be eligible for 10 hours of free credit for Math 106 and 107, provided you get a grade of C, P, or better in Math 208 this semester. To be considered for this benefit, you should register with the Department of Mathematics and Statistics, 810 Oldfather Hall, by Wednesday, February 15,